# Multi-Source Domain Adaptation for Object Detection with Prototype-based Mean Teacher

### v2

@@ 1 Introduction
Object detection (OD), one of the fundamental tasks in computer vision, has made significant progress in recent years [38, 2, 24]. However, this progress, typically measured by performance on curated benchmark datasets such as MS-COCO [16] or Pascal VOC [6], may decline significantly with changes in the data distribution between training (source) and testing (target) domain data [37, 31]. This distribution shift can be caused by several factors, like variations in capture conditions, e.g., weather, illumination, resolution, geographic locations, and size of objects. Since it is too costly to annotate data collected from every operational target domain, researchers have explored a multitude of techniques for unsupervised domain adaptation (UDA), which seeks to find a common representation space between data distributions of a labelled source domain and an unlabeled target domain [3, 22, 15].
@@ 2 Related Works
(c) Prototype-based Learning. This learning paradigm is used in different forms for open-world OD [11], semi-supervised OD [14], domain-adaptive OD [35], and few-shot OD [26]. In open-world OD, prototypes are used to achieve class separation in the feature space and help unknown class identification [11]. In semi-supervised OD, prototypes are used for class distribution alignment between pseudo-labels and their highly overlapping proposals [14]. In the detection of objects across domains, prototypes are used to align the foreground and background regions between the source and target domains [35]. Few-shot OD used universal prototypes to learn the invariant object characteristics across seen and novel classes [26]. In contrast, we use prototypes in MSDA settings to simplify domain-specific feature learning.
@@ 3.2 Learning with Pseudo-labels
$\mathcal{L}_{\mbox{unsup}}=\sum_{i=1}^{M_{T}}\mathcal{L}_{\mbox{cls}}(x_{i},\widetilde{y}_{i})+\mathcal{L}_{\mbox{reg}}(x_{i},\widetilde{y}_{i})$ (2)
where $\widetilde{y}_{i}$ represents the filtered pseudo-labels generated by the teacher model.
@@ 3.4 Domain-specific Features with Prototypes
In our PMT method, prototypes are used to preserve domain-specific information. Preserving domain-specific information helps for distilling with confident pseudo labels from individual source domains. To obtain the prototypes, we added a prototype network to the output of the ROI pool features. This network maps the output of the ROI pool to a $d$ dimensional feature space, where class-conditioned feature alignment is performed using a contrastive loss. This aligns representations of the same object categories across different domains and separates representations of different object categories. Fig. 3 illustrates the prototypes working for domain-specific feature alignment. We produce a local prototype for each class (e.g., car, truck, bus in Fig. 3) in each domain, and these prototypes preserve the domain-specific information. The prototype for each class “across all domains", called global prototypes, is obtained as the mean of the local prototypes of that class. The contrastive loss on global prototypes pushes them apart in the feature space, resulting in better separation of the classes. It also aligns the local prototypes from all domains for each class, reducing confusion due to appearance variation. Note that the prototype network shares parameters across domains, so its parameter size won’t change regardless of the number of domains used.
@@ 4.1 Experimental Methodology
Implementation Details. Our experiments follow the same procedure as in [27, 33]. The FRCNN[21] VGG16 backbone pre-trained with ImageNet was used as the detection framework. Similar to [9], ROI-alignment was used, and the shorter side of the input image was reduced to 600 pixels. For our mean-teacher training, the same setting was used as in [17]. For filtering the pseudo-labels, we set the confidence threshold of 0.7. The model is trained in the burn-in setting for 15 epochs. Then, the model was trained in the unsupervised DA setting for another 15 epochs. For all settings, the value $\alpha$ and $\beta$ were set to 1 and 0.1, respectively, while the value of $\gamma$ was set to 1.2, 0.6, and 0.1 for Sections 4.2, 4.3 and 4.4. The weight smoothing coefficient of EMA is set to 0.9996. All the experiments were conducted on 4 A100 GPUs, with batch size 4 and a learning rate of 0.2. Our approach was implemented using Pytorch[19] and Detectron2[28].
Quantitative comparison. We compared our approach with the following baseline: (1) Source-Only: FRCNN[21] is trained in a supervised manner on source domain data, and no adaptation is performed. (2) UDA Blending: all the source domain data is combined, and UDA methods [22, 32, 29, 17, 15] are used to adapt FRCNN. (3) MSDA: adaptation of FRCNN from independent source datasets. It includes MSDA methods developed for classification [20, 36] and upgraded for OD, as well as MSDA methods developed for OD [33][27]. (4) Oracle: Target-Only and All-Combined. In the Target-Only case, we performed supervised training of FRCNN only on labeled target data. In the All-Combined case, supervised training is performed on all source and target data combined. Note that some of the baseline results are taken from [33] and [27]. Class-wise AP is reported in the supplementary material.
@@ 4.2 Cross Time Adaptation
The mAP performance of our method is compared against the other approaches in Table 1. It can be observed that the UDA methods that blend source datasets are able to increase the performance of the detector compared to the Source-Only baseline. However, this increase in performance is not large, as they disregard the inter-source domain shift. It is also interesting to note that two state-of-the-art MSDA classification methods [36, 20], perform worse than the Source-Only baseline. This degradation in performance shows the fundamental difference in MSDA methods for classification and detection and a strong motivation for MSDA for OD. Our method improves the performance of the best-performing MSDA method [27] by 5.5%. It can also be seen that the performance of our PMT method is much better than the Oracle Target-Only case (because of the small size of the target dataset), and it approached the mAP of the Oracle All-Combined case.
@@ 4.3 Cross Camera Adaptation
Note that in this experiment, there is only one object category (car), so the prototype separation part was removed for our contrastive loss. It can be observed that the variation in the performance is similar to Section 4.2. Our PMT can outperform state-of-the-art methods, but the increase in AP is lower compared to TRKP. We suppose this is due to the removal of the prototype separation part of our contrastive loss. In ablation studies, we showed that every loss component contributed to obtaining the best PMT performance.
@@ 4.4 Extension to Mixed Domain Adaptation
In the MSDA, the domain shift among source data is not always limited to one factor. So, a setting with domain shift with multiple factors is considered for validation. In the source domain, we considered the MS COCO[16], Cityscapes[4], and Synscapes[25] datasets, and the Daytime domain of the BDD100K dataset are the target domain. Among the source domains, the domain shift is mixed and there are more classes, making this study a challenging scenario. Also, the number of source domains is increased from two to three. The source domain has 99,724 (2,975 Cityscapes + 71,749 MS COCO + 25,000 Synscapes) labeled images and, the target domain has 36,728 unlabeled images. The evaluation is performed on 5,258 validation set images of Daytime. For training and evaluation, we only employ the images having 7 common object categories between the datasets. From the results reported in Table 3, we can observe that the Source-Only setting performs better than the UDA Blending setting. This is because of the complex domain shift among the source data. When only Cityscapes and MS COCO are used as the source domain, our method outperforms the previous best-performing method by 3.4%. After adding the Synscapes dataset our method still outperforms other methods, including the Oracle Target-Only results.
@@
4.5 Ablation Study
The cross-time adaptation setting is used in the ablation studies. The two main components of our method are the prototype network and multi-class discriminator. To clearly disentangle the effect of terms in the contrastive loss for prototypes, we considered them as two separate parts: “prototype separation” and “prototype alignment" part. The prototype separation loss increases the distance between different object categories across domains, whereas the prototype alignment loss aligns the same object category across various domains. The experimental results shown in Table 4 show that the results without using the discriminator and prototype are better than the Source-Only result of Table 1. This is due to the effectiveness of the mean-teacher framework in MSDA. By adding the discriminator to the network, performance is increased by 10.2%. This result is already better than the previous state of art MSDA methods for OD. By adding both of our prototype loss terms individually, the performance of the model further increases. Note that using the alignment loss performs better than the separation loss because the separation loss only discriminates the categories, and the vanilla FRCNN is already doing it. Finally, when both the prototype loss terms are combined, the performance is further improved.
@@ 4.5 Ablation Study
$\checkmark$ $\checkmark$ $\checkmark$ 45.3
@@ 4.6 Parameter Growth with Number of Domains
Table 5: The growth of model parameter (in millions) as the number of source domains increases. While the parameters of DMSN [33] and TRKP [27] grow quickly, the parameters of our method PMT remain almost constant because the increase is only due to the prototype vectors.
@@ 5 Conclusion
State-of-the-art approaches for MSDA learn domain-invariant and domain-specific parameters for each source domain to effectively adapt DL models from multiple source domains. In this paper, a new MSDA method is proposed, where domain-specific information is encoded using class prototypes. This avoids the need to train a domain-specific subnet for each source domain, simplifying the MSDA architecture significantly. The number of parameters required for our method remains almost constant regardless of the number of source domains, and it is comparable to adaptation with single-source UDA methods, since only prototype vectors are needed for each source domain. Experimental results indicate that our PMT method can effectively exploit multiple domains and outperform state-of-the-art methods on challenging OD baselines.
Acknowledgement- This work was supported by the MITACS, Natural Sciences and Engineering Research Council of Canada, and the Digital Research Alliance of Canada.

### v3

# Multi-Source Domain Adaptation for Object Detection with Prototype-based Mean-teacher

@@ 1 Introduction
Object detection (OD), one of the fundamental tasks in computer vision, has made significant progress in recent years [40, 2, 26]. However, this progress, typically measured by performance on curated benchmark datasets such as MS-COCO [16] or Pascal VOC [6], may decline significantly with changes in the data distribution between training (source) and testing (target) domain data [39, 33]. This distribution shift can be caused by several factors, like variations in capture conditions, e.g., weather, illumination, resolution, geographic locations, and size of objects. Since it is too costly to annotate data collected from every operational target domain, researchers have explored a multitude of techniques for unsupervised domain adaptation (UDA), which seeks to find a common representation space between data distributions of a labeled source domain and an unlabeled target domain [3, 24, 15].
@@ 2 Related Works
Many DL models have been proposed for OD, namely, Faster-RCNN (FRCNN)[23], Fully Convolutional One Stage (FCOS)[26], You Only Look Once (YOLO)[22], Single Shot Detector (SSD)[17]. Although our PMT method is independent of the OD model, this paper focuses on FRCNN as our base detector for SOTA comparison.
@@
(c) Prototype-based Learning. This learning paradigm is used in different forms for open-world OD [11], semi-supervised OD [14], domain-adaptive OD [37], and few-shot OD [28]. In open-world OD, prototypes are used to achieve class separation in the feature space and help unknown class identification [11]. In semi-supervised OD, prototypes are used for class distribution alignment between pseudo-labels and their highly overlapping proposals [14]. In the detection of objects across domains, prototypes are used to align the foreground and background regions between the source and target domains [37]. Few-shot OD used universal prototypes to learn the invariant object characteristics across seen and novel classes [28]. In contrast, we use prototypes in MSDA settings to simplify domain-specific feature learning. Existing MSDA methods learn domain-specific features by using separate networks for each source domain. We show that by using domain-specific prototype vectors instead, we can avoid learning separate networks for each domain.
@@ 3.2 Learning with Pseudo-labels
$\mathcal{L}_{\mbox{unsup}}=\sum_{i=1}^{M_{T}}\mathcal{L}_{\mbox{cls}}(x_{i},\widetilde{y}_{i})$ (2)
where $\widetilde{y}_{i}$ represents the filtered pseudo-labels generated by the teacher model. Note that the regression loss is not used here, as the confidence thresholding is not sufficient to filter the pseudo-labels that are potentially incorrect for bounding box regression [18].
@@ 3.4 Domain-specific Features with Prototypes
In our PMT method, prototypes are used to preserve domain-specific information. Preserving domain-specific information helps for distilling with confident pseudo labels from individual source domains. To obtain the prototypes, we added a prototype network to the output of the ROI pool features. This network maps the output of the ROI pool to a $d$ dimensional feature space, where class-conditioned feature alignment is performed using a contrastive loss. This aligns representations of the same object categories across different domains and separates representations of different object categories. Fig. 3 illustrates the prototypes working for domain-specific feature alignment. We produce a local prototype for each class (e.g., car, truck, bus in Fig. 3) in each domain, and these prototypes preserve the domain-specific information. The prototype for each class “across all domains”, called global prototypes, is obtained as the mean of the local prototypes of that class. The contrastive loss on global prototypes pushes them apart in the feature space, resulting in better separation of the classes. It also aligns the local prototypes from all domains for each class, reducing confusion due to appearance variation. Note that the prototype network shares parameters across domains, so its parameter size won’t change regardless of the number of domains used.
@@ 4.1 Experimental Methodology
Implementation Details. Our experiments follow the same procedure as in [29, 35]. FRCNN[23] VGG16 backbone pre-trained with ImageNet was used as the detection framework. Similar to [9], ROI-alignment was used, and the shorter side of the input image was reduced to 600 pixels. For our mean-teacher training, the same setting was used as in [18]. For filtering the pseudo-labels a threshold of 0.7 is used. The model is trained in the burn-in setting for 15 epochs. Then, the model was trained in the unsupervised DA setting for another 15 epochs. For all the settings, the value $\alpha$ and $\beta$ were set to 1 and 0.1, respectively, while the value of $\gamma$ was set to 1.2, 0.6, and 0.1 for Sections 4.2, 4.3 and 4.4. The weight smoothing coefficient of EMA is set to 0.9996. We conducted all the experiments on 4 A100 GPUs, with batch size 4 and a learning rate of 0.02. Our approach was implemented using Pytorch[20] and Detectron2[30].
Quantitative comparison. We compared our approach with the following baseline: (1) Source-Only: FRCNN[23] is trained in a supervised manner on source domain data, and no adaptation is performed. (2) UDA Blending: all the source domain data is combined, and UDA methods [24, 34, 31, 18, 15] are used to adapt FRCNN. (3) MSDA: adaptation of FRCNN from independent source datasets. It includes MSDA methods developed for classification [21, 38] and upgraded for OD, as well as MSDA methods developed for OD [35][29]. (4) Oracle: Target-Only and All-Combined. In the Target-Only case, we performed supervised training of FRCNN only on labeled target data. In the All-Combined case, supervised training is performed on all source and target data combined. Note that some of the baseline results are taken from [35] and [29].
@@ 4.2 Cross Time Adaptation
The mAP performance of our method is compared against the other approaches in Table 1. Class-wise AP is reported in Table 7 in Appendix. It can be observed that the UDA methods that blend source datasets are able to increase the performance of the detector compared to the Source-Only baseline. However, this increase in performance is not large, as they disregard the inter-source domain shift. It is also interesting to note that two state-of-the-art MSDA classification methods [38, 21], perform worse than the Source-Only baseline. This degradation in performance shows the fundamental difference in MSDA methods for classification and detection and a strong motivation for MSDA for OD. Our method improves on the performance of the best-performing MSDA method [29] by 5.5%. It can also be seen that the performance of our PMT method is much better than the Oracle Target-Only case (because of the small size of the target dataset), and it approached the mAP of the Oracle All-Combined case.
@@ 4.3 Cross Camera Adaptation
Note that in this experiment, there is only one object category (car), so we cannot use both components of our contrastive loss 9. For this case, the prototype separation part was removed for our contrastive loss. It can be observed that the variation in the performance is similar to Section 4.2. Our PMT can outperform state-of-the-art methods, but the increase in AP is lower compared to TRKP. We suppose this is due to the removal of the prototype separation part of our contrastive loss. In ablation studies, we showed that every loss component contributed to obtaining the best PMT performance.
@@ 4.4 Extension to Mixed Domain Adaptation
In the MSDA, the domain shift among source data is not always limited to one factor. So, a setting with domain shift with multiple factors is considered for validation. In the source domain, we considered the MS COCO[16], Cityscapes[4], and Synscapes[27] datasets, and the Daytime domain of the BDD100K dataset are the target domain. Among the source domains, the domain shift is mixed and there are more classes, making this study a challenging scenario. Also, the number of source domains is increased from two to three. The source domain has 99,724 (2,975 Cityscapes + 71,749 MS COCO + 25,000 Synscapes) labeled images and, the target domain has 36,728 unlabeled images. The evaluation is performed on 5,258 validation set images of Daytime. For training and evaluation, we only employ the images having 7 common object categories between the datasets. From the results reported in Table 3, we can observe that the Source-Only setting performs better than the UDA Blending setting. This is because of the complex domain shift among the source data. When only Cityscapes and MS COCO are used as the source domain, our method outperforms the previous best-performing method by 3.4%. After adding the Synscapes dataset our method still outperforms other methods, including the Oracle Target-Only results. This experiment further highlights the effectiveness of our method. Class-wise AP is reported in Table 8 in Appendix.
@@
4.5 Ablation Studies
Impact of the different components. The cross-time adaptation setting is used in the ablation studies. The two main components of our method are the prototype network and multi-class discriminator. To clearly disentangle the effect of terms in the contrastive loss for prototypes, we considered them as two separate parts: “prototype separation” and “prototype alignment” part. The prototype separation loss increases the distance between different object categories across domains, whereas the prototype alignment loss aligns the same object category across various domains. The experimental results shown in Table 4 show that the results without using the discriminator and prototype are better than the Source-Only result of Table 1. This is due to the effectiveness of the mean-teacher framework in MSDA. By adding the discriminator to the network, performance is increased by 10.2%. This result is already better than the previous state of art MSDA methods for OD. By adding both of our prototype loss terms individually, the performance of the model further increases. Note that using the alignment loss performs better than the separation loss because the separation loss only discriminates the categories, and the vanilla FRCNN is already doing it. Finally, when both the prototype loss terms are combined, the performance is further improved.
@@ 4.5 Ablation Studies
$\checkmark$ $\checkmark$ $\checkmark$ 44.6
@@ 4.6 Parameter Growth with Number of Domains
Table 5: Model parameter growth (in millions) as the number of source domains increases. While the parameters of DMSN [35] and TRKP [29] grow quickly with the number of source domains, the parameters of our method PMT remain almost constant because the increase is only due to the prototype vectors.
@@ 5 Conclusion
A commonly used architecture for MSDA is to learn domain-invariant and domain-specific parameters (for each source domain) to effectively adapt DL models from multiple source domains. In this paper, we proposed a method for simplifying MSDA where domain-specific information is retained using class prototypes for each domain. This avoids the need to train a domain-specific subnet for each source domain, simplifying the MSDA architecture significantly. The parameters of our model remain almost constant regardless of the number of source domains used as only prototype vectors are needed for each domain. Thus, our model size is comparable to UDA methods which perform adaptation from a single source. Experimental results show that our proposed method can effectively exploit multiple domains and improve on the state-of-the-art for multi-source object detection.